OpenStudy (anonymous):

x + 2y = 10 2x +3y = 16

OpenStudy (anonymous):

you need to eliminate one of the variables

OpenStudy (anonymous):

i know but i need help working it out

OpenStudy (anonymous):

subtract 2y from the left side of the first equation so it becomes x=-2y+10

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

no like show me how like draw it out

OpenStudy (anonymous):

Okay so that turns into x=-2y+10 2x+3y=16

OpenStudy (anonymous):

Now you plug in the first equation for x in the second one like so: 2(-2y+10)+3y=16

OpenStudy (anonymous):

Now distribute.

OpenStudy (anonymous):

Y=4

OpenStudy (anonymous):

Not quite. -4y+20+3y=16 -y+20=16 -y=-36 y=36

OpenStudy (anonymous):

cause when you distrubte it is -4y=20=3y+16

OpenStudy (anonymous):

Not exactly. Look at my last post.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

So now that you know y=36, plug it into the first equation.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

x+2(36)=10 x+72=10 x=-62 So your coordinates are (-62,36)